OpenStudy (anonymous):

840_10= the under score

OpenStudy (anonymous):

840\[-\] 10

jimthompson5910 (jim_thompson5910):

When dividing by 10, simply move the decimal point to the left one spot So to divide 840 by 10, think of 840 as 840.0 and move the decimal point to the left one spot to get 84.0 which is 84 So \[\large \frac{840}{10}=84\]

OpenStudy (anonymous):

Thats the line i wanted thank you!! So you just divide it :D

jimthompson5910 (jim_thompson5910):

alternatively, you can cancel the trailing zeros. So the zero in 840 and the zero in 10 cancel out leaving you with \[\large \frac{840}{10}=\frac{84}{1}=84\]

jimthompson5910 (jim_thompson5910):

yes fractions are another way of saying divide

jimthompson5910 (jim_thompson5910):

btw you represent the problem as 840/10 instead of 840_10

OpenStudy (anonymous):

Omg Thank you!! You mkae everything so understandable